OpenStudy (anonymous):

Write an equation with a variable on both sides of the equal sign that has infinitely many solutions. Solve the equation and explain why it has an infinite number of solutions.

OpenStudy (anonymous):

@Butterfly16 can you help me?

OpenStudy (anonymous):

Hmm, try this one :o) 10x-2=2(x+15)

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Can you find the solution? :o)

OpenStudy (anonymous):

iniqualitys confuse me how do you get the solution?

OpenStudy (anonymous):

Okay, first, lets solve for x, so do you know our first step?

OpenStudy (anonymous):

im back um im pretty sure you deal with the perenthesis first

OpenStudy (anonymous):

Yup, first we distribute: \[10x-2=2(x+15) \rightarrow 10x-2=2x+30\]